# Only the Ambidextrous Can Flock: Two-dimensional Chiral Malthusian Flocks, Time crystals, and the KPZ Equation

### v1

@@ Abstract
We study two-dimensional chiral dry Malthusian flocks; that is, chiral polar-ordered active matter with neither number nor momentum conservation. In the absence of fluctuations, these form a “time crystal”, in which the velocity rotates uniformly at a fixed frequency. Fluctuations are described by the ($2+1$)-Kardar-Parisi-Zhang (KPZ) equation, which implies short-ranged orientational order. For weak chirality, the system is in the linear regime of the KPZ equation for a wide range of length scales, over which it exhibits quasi-long-ranged orientational order. Our predictions for velocity and density correlations are testable in both simulations and experiments.
@@
Figure 1: (a) In a generic 2D chiral Malthusian flock, the mean velocity of the whole flock is denoted by $\bar{\mathbf{v}}$ (red) , where $\bar{\mathbf{v}}=v_{0}[\cos(bt)\hat{\bf x}-\sin(bt)\hat{\bf y}]$, while the spatio-temporally fluctuating velocity field is denoted by $\mathbf{v}$ (blue) where $\mathbf{v}=v_{0}[\cos(bt+\phi)\hat{\bf x}-\sin(bt+\phi)\hat{\bf y}]$. Ignoring variations in $|\mathbf{v}|$, all fluctuations can be captured by the angular field $\phi(x,y,t)$. (b) Here, we show that the hydrodynamic behavior of $\phi$ can generically be mapped on the Kardar-Parisi-Zhang surface growth model [30] in (2+1) dimensions. [The image in (b) is reproduced with permission from EPL 109, 46003 (2015) [42].
@@
In our discussion of low chirality systems in this paper, we will assume that the chirality is lowered by “racemic mixing”, so that all chiral coefficients vanish at the same concentration. Other possibilities will be discussed in [23].
@@
Of course, in a time-averaged sense, rotation invariance is not broken in a state like (1); all directions of velocity are sampled for equal time. The symmetry that is broken by (1) is time-translation invariance, which is broken because the system spontaneously chooses a particular phase $\phi$ in (1), which is equivalent to choosing a particular set of discrete times for the velocity to point, say, along $\bf{\hat{x}}$. In this sense, the state (1) can be thought of as a type of “time crystal”: a system that is periodic in time in the same way that a conventional crystal is periodic in space.
@@
$\mathbf{v}(\mathbf{r},t)=v_{0}\left[\cos(bt+\phi(\mathbf{r},t))\hat{\mathbf{x}}-\sin(bt+\phi(\mathbf{r},t))\hat{\mathbf{y}}\right]$ (5)
@@
$\partial_{t}v_{i}=(b+\partial_{t}\phi)\epsilon_{ik}v_{k}\ \ \ ,\ \ \ \partial_{j}v_{i}=\epsilon_{ik}v_{k}\partial_{j}\phi{\color[rgb]{0,0,1}\,.}$ (6)
@@
Second, we time-average Eq. (8) over one oscillation cycle of $\mathbf{v}$. For the purposes of such time averages, $\phi(\mathbf{r},t)$ and all of its spatiotemporal derivatives can be treated as constants, since $\phi(\mathbf{r},t)$ varies slowly on the time scale $b^{-1}$ of the oscillating velocity. They can therefore be removed from the averages over one cycle, which we’ll denote by $\langle\rangle_{c}$. It is then fairly straightforward to show that time averages of products of two and four components of $\mathbf{v}$ are given by:
@@
$\partial_{t}\phi=\nu\nabla^{2}\phi+{\lambda_{K}\over 2}(\nabla\phi)^{2}+f_{\phi}\,,$ (11)
where $\nu=\mu_{1}+\left({\mu_{2}+\mu_{3}v_{0}^{2}\over 2}\right)$, $\lambda_{K}=-\mu^{\prime}$ and where $f_{\phi}$ is a Gaussian, zero-mean white noise with the following statistics:
@@
We note that the KPZ universality class (UC) encompasses an amazing array of diverse systems–recent additions to this include incompressible polar active fluids [26], driven Bose-Einstein condensates [27], and non-reciprocal systems [28, 29]; It is nonetheless surprising, in our opinion, that all generic 2D chiral Malthusian flocks necessarily fall into the KPZ UC as well.
@@
Small chirality limit—When the chirality $b$ is small, the system looks, over a large range of length and time scales, as if it were achiral. Over these length and time scales, a weakly chiral system will exhibit “anomalous hydrodynamics”, with its effective damping coefficients growing with length scale. This leads, as we show in detail in the (ALP) [23], to a strong dependence of the parameters $\nu$ and $\lambda_{K}$ in our KPZ equation on the chirality $b$; they respectively diverge and vanish in the limit of small $b$; for the “racemic” case described above
$\nu(b)\propto b^{-\eta_{\nu}}\ \ \ ,\ \ \ \lambda_{K}(b)\propto b\,,$ (17)
@@
The numerical value given here is based on numerical evaluation of the noisy hydrodynamic equation for achiral Malthusian flocks by Ref. [31].
@@
Figure 2: Regimes of different behavior in the limit of weak chirality. The longest length scales $L_{NL}$ and $L_{v}$ and time scales $\tau_{{}_{NL}}$ and $\tau_{v}$ will be long enough to allow several decades of length and time scale to lie in each region, making our scaling predictions for each region experimentally accessible. See the text for more details.
@@
$L_{{}_{NL}}(b)$ $\propto$ $\exp\left(C_{{}_{NL}}b^{-\eta_{g}}\right)\ \ \ ,\ \ \ \eta_{g}\approx 19/5\,,$ (20)
$L_{v}(b)$ $\propto$ $\exp\left(C_{vL}b^{-\eta_{y}}\right)\ \ \ ,\ \ \ \eta_{y}\approx 22/5\,,$ (21)
@@
where $C_{{}_{NL}}$, $C_{vL}$, and $C_{v\tau}$ are all non-universal constants. The exponents $z_{a}$, $\eta_{g}$, and $\eta_{y}$ are all related to the anomalous hydrodynamic exponents characterizing the scaling of achiral flocks, whose numerical values shown are again based on the numerical results of Ref. [31].
@@
Linear KPZ regime—Here, the effects of the $\lambda_{K}$ non-linearity in the KPZ equation are negligible. Dropping that term leaves an analytic solvable linear theory, from which we predict that the velocity correlations oscillate rapidly in time and decay algebraically in space and time:
$\langle\mathbf{v}(\mathbf{r},t)\cdot\mathbf{v}(\mathbf{r}^{\prime},t^{\prime})\rangle\approx v_{0}^{2}\cos\left[b(t-t^{\prime})\right]\times\left\{\begin{array}[]{ll}\sqrt{{\rm e}^{\alpha\gamma}}\left(|\mathbf{r}-\mathbf{r}^{\prime}|\over L_{c}\right)^{-\alpha}\ \ \ ,&\left(\frac{|\mathbf{r}-\mathbf{r}^{\prime}|}{L_{c}}\right)^{2}\gg\frac{|t-t^{\prime}|}{\tau_{c}}\ \ \ ,\ \ \ \left(\frac{|\mathbf{r}-\mathbf{r}^{\prime}|}{L_{c}}\right)^{2}\gg 1\,,\\ \left(|t-t^{\prime}|\over\tau_{c}\right)^{-{\alpha\over 2}}\ \ \ ,&\frac{|t-t^{\prime}|}{\tau_{c}}\gg\left(\frac{|\mathbf{r}-\mathbf{r}^{\prime}|}{L_{c}}\right)^{2}\ \ \ ,\ \ \ \frac{|t-t^{\prime}|}{\tau_{c}}\gg 1\,,\end{array}\right.\,$ (27)
@@
$\alpha={D_{\phi}\over 2\pi\nu}=C_{\alpha}b^{\eta_{\nu}\over z_{a}}\ \ \ ,\ \ \ \eta_{\nu}\approx{3\over 5}\ \ \ ,\ \ \ z_{a}\approx{5\over 4}\ \ \ ,$ (28)
@@
Note that for small chirality $b$, the exponent $\alpha$ will be small. Hence, the algebraic decay of velocity correlations (27) will be extremely slow; this is what we meant by our earlier cryptic comment that “some regimes are effectively ordered”. Strictly speaking, the order in this regime is “quasi-long-ranged”.
@@
Nonlinear KPZ regime— The effects of the $\lambda_{K}$ non-linearity in the KPZ equation now dominate. The mean squared difference between the phases at widely separated spatial and temporal positions $(\mathbf{r},t)$ and $(\mathbf{r}^{\prime},t^{\prime})$ are given by (15).
If the fluctuations of the field $\phi$ were Gaussian, then (15) would imply that the equal-time correlation of the velocity would be stretched exponentially decaying velocity correlations. However, since the $\phi$ fluctuations are non-Gaussian, due to the relevant nonlinearity $\lambda_{K}$ in the KPZ equation, we can say very little, beyond noting that we expect velocity correlations to decay rapidly with increasing separation $|\mathbf{r}-\mathbf{r}^{\prime}|$ once that separation is larger than $L_{NL}$.
@@
is also interesting in this regime (and experimentally measurable). Unlike the equal-time case, the equal-position velocity correlations carry frequencies $\omega=\pm b$ (see the cosine factor in (27)). This is reminiscent of the spatially periodic density modulation in crystals. Just as that spatial modulation leads to Bragg peaks in the X-ray scattering, the temporal periodicity of our system leads to “Bragg peaks” in the temporal Fourier transform of the velocity correlation at frequencies $\omega=\pm b$ in frequency:
@@
In summary, we have shown that chiral Malthusian flocks in two dimensions belong generically to the KPZ universality. As a result, chirality always disorders a 2D chiral Malthusian flock, leading to short-ranged velocity correlations. We have further shown that for small chirality, the velocity can remain well correlated out to quite long distances. Finally, we have made a number of very detailed predictions for the behavior of these correlation functions that are testable in both simulations and experiments on chiral systems.

### v2

# Only the Ambidextrous Can Flock: Two-dimensional Chiral Malthusian Flocks, Time Cholesterics, and the KPZ Equation

@@ Abstract
We study two-dimensional chiral dry Malthusian flocks; that is, chiral polar-ordered active matter with neither number nor momentum conservation. In the absence of fluctuations, these form a “time cholesteric”, in which the velocity rotates uniformly in time at a fixed frequency. Fluctuations are described by the ($2+1$)-Kardar-Parisi-Zhang (KPZ) equation, which implies short-ranged orientational order. For weak chirality, the system is in the linear regime of the KPZ equation for a wide range of length scales, over which it exhibits quasi-long-ranged orientational order. Our predictions for velocity and density correlations are testable in both simulations and experiments.
@@
Figure 1: (a) In a generic 2D chiral Malthusian flock, the mean velocity of the whole flock is denoted by $\bar{\mathbf{v}}$ (red) , where $\bar{\mathbf{v}}=v_{0}[\cos(bt)\hat{\bf x}-\sin(bt)\hat{\bf y}]$, while the spatio-temporally fluctuating velocity field is denoted by $\mathbf{v}$ (blue) where $\mathbf{v}=v_{0}[\cos(bt+\phi)\hat{\bf x}-\sin(bt+\phi)\hat{\bf y}]$. Ignoring variations in $|\mathbf{v}|$, all fluctuations can be captured by the angular field $\phi(x,y,t)$. (b) Here, we show that the hydrodynamic behavior of $\phi$ can generically be mapped on the Kardar-Parisi-Zhang surface growth model [25] in (2+1) dimensions. The image in (b) is reproduced with permission from EPL 109, 46003 (2015) [26].
@@
In our discussion of low chirality systems in this paper, we will assume that the chirality is lowered by “racemic mixing” of chiral components that are perfect mirror images of each other. In such systems, all chiral coefficients vanish at the same concentration. Other possibilities will be discussed in [23].
@@
Of course, in a time-averaged sense, rotation invariance is not broken in a state like (1); all directions of velocity are sampled for equal time. The symmetry that is broken by (1) is time-translation invariance, which is broken because the system spontaneously chooses a particular phase $\phi$ in (1), which is equivalent to choosing a particular set of discrete times for the velocity to point, say, along $\bf{\hat{x}}$. In this sense, the state (1) can be thought of as a type of “time cholesteric”: the velocity rotates in time in the same way that the molecular axis rotates in space in a cholesteric liquid crystal. That is, the active chiral flock can thought of as a cholesteric with the “pitch axis” being time[24].
@@
$\mathbf{v}(\mathbf{r},t)=v_{0}\left[\cos(bt+\phi(\mathbf{r},t))\hat{\mathbf{x}}-\sin(bt+\phi(\mathbf{r},t))\hat{\mathbf{y}}\right]\,.$ (5)
@@
$\partial_{t}v_{i}=(b+\partial_{t}\phi)\epsilon_{ik}v_{k}\ \ \ ,\ \ \ \partial_{j}v_{i}=\epsilon_{ik}v_{k}\partial_{j}\phi\,.$ (6)
@@
Second, we time-average Eq. (8) over one oscillation cycle of $\mathbf{v}$. For the purposes of such time averages, $\phi(\mathbf{r},t)$ and all of its spatio-temporal derivatives can be treated as constants, since $\phi(\mathbf{r},t)$ varies slowly on the time scale $b^{-1}$ of the oscillating velocity. They can therefore be removed from the averages over one cycle, which we’ll denote by $\langle\rangle_{c}$. It is then fairly straightforward to show that time averages of products of two and four components of $\mathbf{v}$ are given by:
@@
$\partial_{t}\phi=\nu\nabla^{2}\phi+{\lambda_{{}_{K}}\over 2}(\nabla\phi)^{2}+f_{\phi}\,,$ (11)
where $\nu=\mu_{1}+\left({\mu_{2}+\mu_{3}v_{0}^{2}\over 2}\right)$, $\lambda_{{}_{K}}=-\mu^{\prime}$ and where $f_{\phi}$ is a Gaussian, zero-mean white noise with the following statistics:
@@
We note that the KPZ universality class (UC) encompasses an amazing array of diverse systems–recent additions to this include incompressible polar active fluids [29], driven Bose-Einstein condensates [30], and non-reciprocal systems [31, 32]. It is nonetheless surprising, in our opinion, that all generic 2D chiral Malthusian flocks necessarily fall into the KPZ UC as well.
@@
Small chirality limit—When the chirality $b$ is small, the system looks, over a large range of length and time scales, as if it were achiral. Over these length and time scales, a weakly chiral system will exhibit “anomalous hydrodynamics”, with its effective damping coefficients growing with length scale. This leads, as we show in detail in the (ALP) [23], to a strong dependence of the parameters $\nu$ and $\lambda_{{}_{K}}$ in our KPZ equation on the chirality $b$; they respectively diverge and vanish in the limit of small $b$; for the “racemic” case described above
$\nu(b)\propto b^{-\eta_{\nu}}\ \ \ ,\ \ \ \lambda_{{}_{K}}(b)\propto b\,,$ (17)
@@
The numerical value given here is based on numerical simulations of the noisy hydrodynamic equation for achiral Malthusian flocks by Ref. [33].
@@
Figure 2: Regimes of different behavior in the limit of weak chirality. The longest length scales $L_{{}_{NL}}$ and $L_{v}$ and time scales $\tau_{{}_{NL}}$ and $\tau_{v}$ will be long enough to allow several decades of length and time scale to lie in each region, making our scaling predictions for each region experimentally accessible. See the text for more details.
@@
$L_{{}_{NL}}(b)$ $\propto\exp\left(C_{{}_{NL}}b^{-\eta_{g}}\right)\ ,\ \eta_{g}=6/z_{a}-1\approx 19/5\,,$ (20)
$L_{v}(b)$ $\propto\exp\left(C_{vL}b^{-\eta_{y}}\right)\ ,\ \eta_{y}=8/z_{a}-2\approx 22/5\,,$ (21)
@@
where $C_{{}_{NL}}$, $C_{vL}$, and $C_{v\tau}$ are all non-universal constants.
@@
Linear KPZ regime—Here, the effects of the $\lambda_{{}_{K}}$ non-linearity in the KPZ equation are negligible. Dropping that term leaves an analytic solvable linear theory, from which we predict that the velocity correlations oscillate rapidly in time and decay algebraically in space and time:
$\langle\mathbf{v}(\mathbf{r},t)\cdot\mathbf{v}(\mathbf{r}^{\prime},t^{\prime})\rangle\approx v_{0}^{2}\cos\left[b(t-t^{\prime})\right]\times\left\{\begin{array}[]{ll}\sqrt{{\rm e}^{-\alpha\gamma}}\left(|\mathbf{r}-\mathbf{r}^{\prime}|\over 2L_{c}\right)^{-\alpha}\ \ \ ,&\left(\frac{|\mathbf{r}-\mathbf{r}^{\prime}|}{L_{c}}\right)^{2}\gg\frac{|t-t^{\prime}|}{\tau_{c}}\ \ \ ,\ \ \ \frac{|\mathbf{r}-\mathbf{r}^{\prime}|}{L_{c}}\gg 1\,,\\ \left(|t-t^{\prime}|\over\tau_{c}\right)^{-{\alpha\over 2}}\ \ \ ,&\frac{|t-t^{\prime}|}{\tau_{c}}\gg\left(\frac{|\mathbf{r}-\mathbf{r}^{\prime}|}{L_{c}}\right)^{2}\ \ \ ,\ \ \ \frac{|t-t^{\prime}|}{\tau_{c}}\gg 1\,,\end{array}\right.\,$ (27)
@@
$\alpha={D_{\phi}\over 2\pi\nu}=C_{\alpha}b^{\eta_{\nu}}\ \ \ ,\ \ \ \eta_{\nu}\approx{3\over 5}\ ,$ (28)
@@
Note that for small chirality $b$, the exponent $\alpha$ will be small, since $\nu(b)$ gets very large in that limit. Hence, the algebraic decay of the equal-time velocity correlations obtained by setting $t=t^{\prime}$ in (27) will be extremely slow; this is what we meant by our earlier cryptic comment that “some regimes are effectively ordered”. Strictly speaking, the order in this regime is “quasi-long-ranged”.
@@
is also interesting in this regime (and experimentally measurable). Unlike the equal-time case, the non-equal-time velocity correlations carry frequencies $\omega=\pm b$ (see the cosine factor in (27)). This is reminiscent of the spatially periodic density modulation in crystals. Just as that spatial modulation leads to Bragg peaks in the X-ray scattering, the temporal periodicity of our system leads to “Bragg peaks” in the temporal Fourier transform of the velocity correlation at frequencies $\omega=\pm b$:
@@
Nonlinear KPZ regime— The effects of the $\lambda_{{}_{K}}$ non-linearity in the KPZ equation now dominate. The mean squared difference between the phases at widely separated spatial and temporal positions $(\mathbf{r},t)$ and $(\mathbf{r}^{\prime},t^{\prime})$ are given by (15).
If the fluctuations of the field $\phi$ were Gaussian, then (15) would imply that the equal-time correlation of the velocity would be stretched exponentially decaying velocity correlations. However, since the $\phi$ fluctuations are non-Gaussian, due to the relevant nonlinearity $\lambda_{{}_{K}}$ in the KPZ equation, we can say very little, beyond noting that we expect velocity correlations to decay rapidly with increasing separation $|\mathbf{r}-\mathbf{r}^{\prime}|$ once that separation is larger than $L_{{}_{NL}}$.
In summary, we have shown that chiral Malthusian flocks in two dimensions belong generically to the KPZ universality. As a result, chirality always disorders a 2D chiral Malthusian flock, leading to short-ranged velocity correlations. We have further shown that for small chirality, the velocity can remain well correlated out to quite long distances. Finally, we have made a number of very detailed predictions for the behavior of these correlation functions that are testable in both simulations and experiments on active chiral systems.
@@
We also thank Tim Halpin-Healy and Mehran Kardar for useful discussions about the current state of our understanding of the KPZ equation, and for calling references [37, 38] and [39] to our attention. Last but not least, we are grateful to Ananyo Maitra for valuable discussions about prior work on chirality in Malthusian flocks, and for suggesting the name “time cholesteric”.
Note added - After this work was completed, we learned that the conclusion that chiral Malthusian flocks are described by the KPZ equation had been independently reached by Ananyo Maitra[45], whom we thank for calling this to our attention.